7.1 Madsen Motors's bonds have 9 years remaining to maturity. Interest is paid annually, they have a $1,000 par value, t...

7.1

Madsen Motors's bonds have 9 years remaining to maturity. Interest is paid annually, they have a $1,000 par value, the coupon interest rate is 11%, and the yield to maturity is 13%. What is the bond's current market price? Round your answer to the nearest cent.
Answer #1

Annual coupon=1000*11%=110

Hence current price=Annual coupon*Present value of annuity factor(13%,9)+$1000*Present value of discounting factor(13%,9)

=110*5.131655128+$1000*0.332884833

=$897.37(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=110[1-(1.13)^-9]/0.13

=110*5.131655128

2.Present value of discounting factor=1000/1.13^9

=1000*0.332884833